# Deep Metric Imitation Learning for Stable Motion Primitives

### v2

@@ B.1 Comments on local stability
Considering a one-dimensional sphere, it is notable that at $\alpha=\pi$ (with respect to the goal), both the great-circle distance and the chordal distance can decrease in two possible ways: by evolving either to the right or to the left (see Fig. 15). Consequently, to ensure these distances decrease in the region around $\alpha=\pi$, one of these options should be selected. However, this would require the dynamical system to instantly change its direction at this point, rendering $f^{\mathcal{T}}_{\theta}$ discontinuous. Yet, in Theo. 4, we assume this function is continuous, as continuity is a prerequisite for the class-$\mathcal{KL}$ upper bound $\beta$ to also be continuous, which is necessary to prove stability.

### v3

@@ B.1 Comments on local stability
Considering a one-dimensional sphere, it is notable that at $\alpha=\pi$ (with respect to the goal), both the great-circle distance and the chordal distance can decrease in two possible ways: by evolving either to the right or to the left (see Fig. 15). Consequently, to ensure these distances decrease in the region around $\alpha=\pi$, one of these options should be selected. However, this would require the dynamical system to instantly change its direction at this point, rendering $f^{\mathcal{T}}_{\theta}$ discontinuous. Yet, in Theorem 4, we assume this function is continuous, as continuity is a prerequisite for the class-$\mathcal{KL}$ upper bound $\beta$ to also be continuous, which is necessary to prove stability.